OpenStudy (anonymous):

The Royals softball team played 75 games and won 55 of them. What percent of the games did they lose? (round to the nearest tenth??

OpenStudy (compassionate):

Hi, So they played 75 games (100%) and only won 55. Simply divide 55 into 75. \[\frac{ 75 }{ 55 }\]

OpenStudy (michele_laino):

please note that the games lost are 75-55=20, so if we apply the standard definition of probability, namely favorable cases over possible cases, we get?

OpenStudy (anonymous):

They played 75 games and only won 55 of them: 55/75 Simplify it to get: 11/15 Its asking you to find the percent of the games they lost so, you subtract 15-11=4 4/15=x/100 Proportion You'll get 26.7% if you round your answer to the nearest tenth.

OpenStudy (mathstudent55):

They played 75 games and won 55. The number of games they lost is 75 - 55 = 20 The percentage of games they lost is: \(\dfrac{20}{75} \times 100\% = 26.66666...\%\) Since they ask to round to the nearest tenth, the answer is: They lost 26.7% of the games.

OpenStudy (triciaal):

@compassionate please read the question before answering. in addition, the percentage is usually a fraction of the whole so will be less than 100 the whole = amount loss + amount won

OpenStudy (anonymous):

How many ties?

OpenStudy (anonymous):

zero ties

OpenStudy (anonymous):

20/75*100= 26.7 %

OpenStudy (anonymous):

55 divided by 75 that how you get the percentage. Then time the number by one hundred then round up or down. 55 divided by 75 equals 73%